\begin{thebibliography}{99}
\bibitem{david-ralf} D.\ Asper\'o and R.\ Schindler, {\em Bounded Martin's Maximum with an asterisk}, Notre Dame J.\ Formal Logic 55 (2014), pp.\ 333-348.
\bibitem{cs} B.\ Claverie and R.\ Schindler, {\em Increasing $u_2$ by a stationary set preserving forcing}, Journal of Symb.\ Logic 74 (2009), pp.\ 187-200.
\bibitem{cohen} P.\ Cohen, {\em Set theory and the continuum hypothesis}, Benjamin, New York, 1966.
\bibitem{new?} S.\ Feferman, H.\ Friedman, P.\ Maddy, and J.\ Steel, {\em Does mathematics need new axioms?}, Bulletin of Symbolic Logic \textbf{6} (2000), pp.\ 401-446.  
\bibitem{FMW} Q.\ Feng, M.\ Magidor, and W.H.\ Woodin H., {\em Universally Baire Sets of Reals}, in: Judah H., Just W., Woodin H. (eds) Set Theory of the Continuum. Mathematical Sciences Research Institute Publications, vol 26, Springer, New York, NY 1992.
\bibitem{FMS} M.\ Foreman, M.\ Magidor, and S.\ Shelah, {\em Martin's Maximum, saturated ideals and non-regular ultrafilters I}, Ann.\ of Mathematics 127 (1988), pp.\
1 - 47.
\bibitem{hilbert} D.\ Hilbert, {\em Mathematische Probleme}, Nachr.\ K.\ Ges.\ Wiss.\ G\"ottingen, Math.-Phys.\ Klasse (G\"ottinger Nachrichten), \textbf{3} (1900), pp.\ 253-297. 
\bibitem{stacking} R.\ Jensen, E.\ Schimmerling, R.\ Schindler, and J.\ Steel, {\em Stacking mice}, Journal of Symb.\ Logic 74 (2009), pp.315-335.
\bibitem{peter} P.\ Koellner, {\em The Continuum Hypothesis}, The Stanford Encyclopedia of Philosophy (Spring 2019 Edition), E.\ N.\ Zalta (ed.),  {\tt https://plato.stanford.edu/archives/spr2019/entries/continuum-hypothesis/ }
\bibitem{plarson} P.\ Larson, {\em Martin's Maximum and the Pmax axiom $(*)$}, Annals of Pure and Applied Logic 106 (2000), pp.\ 135-149.
\bibitem{plarson2} P.\ Larson, {\em Martin's Maximum and definability in $H(\omega_2)$}, Annals of Pure and Applied Logic 156 (2008), pp.\ 110-122. 
\bibitem{menachem} M.\ Magidor, {\em Some set theories are more equal},
{\tt http://logic.harvard.edu/EFI${}_-$Magidor.pdf}
\bibitem{proj-determinacy} D.A.\ Martin and J.\ Steel, {\em A Proof of Projective Determinacy}, Journal of the American Mathematical Society 2, pp.\ 71-125. 
\bibitem{moore} J.\ Moore, {\em What makes the continuum $\aleph_2$}, in: Foundations of mathematics, essays in honor of W.\ Hugh Woodin's 60th birthday, Harvard University (Caicedo et al., eds.), pp.\ 259-287. 
\bibitem{trang-sargsyan} G.\ Sargsyan and N.\ Trang, {\em The largest Suslin Axiom},
available at {\tt https://www.math.uci.edu/$\sim$ntrang/lsa.pdf}
\bibitem{both} R.\ Schindler, {\em Woodin's axiom $(*)$, or Martin's Maximum, or both?}, in: Foundations of mathematics, essays in honor of W.\ Hugh Woodin's 60th birthday, Harvard University (Caicedo et al., eds.), pp.\ 177-204.
\bibitem{john} J.\ Steel, {\em {\sf PFA} implies {\sf AD}${}^{L({\mathbb R})}$},
J.\ Symb.\ Logic 70 (2005), pp.\ 1255-1296.
\bibitem{forcing-free} J.\ Steel, {\em A stationary-tower-free proof of the derived model theorem}, in: ``Advances of logic,'' Contemporary Mathematics 425, S.\ Gao, S.\ Jackson, and Y.\ Yang (eds.), Americ.\ Math.\ Society, Providence RI (2007), pp.\ 105-193.
\bibitem{stevo} S.\ Todorcevic, {\em The power set of $\omega_1$ and the Continuum Problem}, {\tt http://logic.harvard.edu/Todorcevic${}_-$Structure4.pdf}
\bibitem{hugh} W.\ H.\ Woodin, {\em The axiom of determinacy, forcing axioms, and
the non-stationary ideal}, de Gruyter, Berlin-New York 1999.
\bibitem{hugh-ch1} W.\ H.\ Woodin, {\em The Continuum Hypothesis, Part I}, Notices 
Amer.\ Math.\ Soc.\ \textbf{48 (6)} (2001), pp.\ 567-576.
\bibitem{hugh-ch2} W.\ H.\ Woodin, {\em The Continuum Hypothesis, Part II}, Notices 
Amer.\ Math.\ Soc.\ \textbf{48 (7)} (2001), pp.\ 681-690.
\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{fancyhdr, array,calc,graphicx,url,tabularx}
\usepackage{geometry}
\usepackage{latexsym}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{makeidx}
\usepackage{cleveref}
\usepackage{enumerate}
\usepackage{verbatim}
\usepackage{tikz}
\usepackage[colorlinks=true,linkcolor=blue]{hyperref}
\allowdisplaybreaks

\usepackage{changepage}

\makeindex

{
   \end{minipage}
   \vspace*{\stretch{3}}
   \clearpage
}

\renewcommand{\models}{\vDash}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{claim}[theorem]{Claim}

\numberwithin{figure}{section}

\newenvironment{Claim}{\noindent{\it
Claim} {}}{}
\newenvironment{Case}{\noindent{\it
Case}}{}

\def\and{\mathrel{\kern1pt\&\kern1pt}}

\def\<#1>{\langle\,#1\,\rangle}

 \input xy
 \xyoption{all}

\title{${\sf MM}^{++}$ implies $(*)$}

\author{David Asper\'o\footnote{Funded by EPSRC Grant EP/N032160/1.} \ and Ralf Schindler\footnote{Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under
Germany’s Excellence Strategy EXC 2044 390685587, Mathematics M\"unster: Dynamics - Geometry
- Structure.}}   
         
\date{\today}

\begin{document}
   
\maketitle

\begin{abstract}
We show that Martin's Maximum${}^{++}$ implies Woodin's ${\mathbb P}_{\rm max}$ axiom $(*)$.
\end{abstract}

\section{Introduction.}

Cantor's Continuum Problem, which later became Hilbert's first Problem (see \cite{hilbert}), asks how many real numbers there are. This question got a non-answer through the discovery of the method of {\em forcing} by Paul Cohen: {\sf CH}, Cantor's Continuum Hypothesis, is independent from {\sf ZFC} (see \cite{cohen}), the standard axiom system for set theory which had been isolated by Zermelo and Fraenkel. {\sf CH} states that every uncountable set of reals has the same size as ${\mathbb R}$.

Ever since Cohen, set theorists have been searching for natural new axioms which extend {\sf ZFC} and which settle the Continuum Problem. See e.g.\ \cite{hugh-ch1}, \cite{hugh-ch2}, \cite{peter}, and the discussion in
\cite{new?}. There are two prominent such axioms which decide {\sf CH} in the negative and which in fact both prove that there are $\aleph_2$ reals: Martin's Maximum ({\sf MM}, for short) or variants thereof on the one hand (see \cite{FMS}), and Woodin's axiom $(*)$ on the other hand (see \cite{hugh}). See e.g.\ \cite{menachem}, \cite{stevo}, and \cite{moore}.

Both of these axioms may be construed as maximality principles for the theory of the structure $(H_{\omega_2};\in)$, but up to this point the relationship between {\sf MM} and $(*)$ was a bit of a mystery, which led M.\ Magidor to call $(*)$ a ``competitor'' of {\sf MM} (\cite[p.\ 18]{menachem}). Both {\sf MM} and $(*)$ are inspired by and formulated in the language of forcing, and they both have ``the same intuitive motivation: Namely, the
universe of sets is rich'' (\cite[p.\ 18]{menachem}). 

This paper resolves the tension between {\sf MM} and $(*)$ by proving that ${\sf MM}^{++}$, a strengthening of {\sf MM}, actually {\em implies} $(*)$, see Theorem \ref{main_thm} below, so that {\sf MM} and $(*)$ are actually compatible with each other.
This answers \cite[Question (18) a) on p.\ 924]{hugh}, see also \cite[p.\ 846]{hugh}, \cite[Conjecture 6.8 on p.\ 19]{menachem},
and \cite[Problem 14.7]{moore}.

\section{Preliminaries.}

{\em Martin's Maximum}${}^{++}$, abbreviated by {\sf MM}${}^{++}$, see \cite{FMS} (cf.\ also \cite[Definition 2.45 (2)]{hugh}), is the statement that if ${\mathbb P}$ is a forcing which preserves stationary subsets of $\omega_1$, if $\{ D_i \colon
i<\omega_1 \}$ is a collection of dense subsets of ${\mathbb P}$, and if $\{ \tau_i
\colon i<\omega_1 \}$ is a collection of ${\mathbb P}$-names for stationary subsets
of $\omega_1$, then there is a filter $g \subset {\mathbb P}$ such that 
for every $i<\omega_1$,
\begin{enumerate}
\item[(i)]
$g \cap
D_i \not= \emptyset$ and 
\item[(ii)] $(\tau_i)^g = \{ \xi<\omega_1
\colon \exists p \in g \, p \Vdash_{\mathbb P} {\check \xi} \in \tau_i \}$ is stationary.
\end{enumerate}

Woodin's ${\mathbb P}_{\rm max}$ axiom (*), see \cite[Definition 5.1]{hugh},
is the statement that 
\begin{enumerate}
\item[(i)]
{\sf AD} holds in $L({\mathbb R})$ and 
\item[(ii)] there is some $g$
which is ${\mathbb P}_{\rm max}$-generic over $L({\mathbb R})$ such that
${\cal P}(\omega_1) \cap V \subset L({\mathbb R})[g]$.
\end{enumerate}

Already {\sf PFA}, the {\em Proper Forcing Axiom}, which is weaker than {\sf MM}${}^{++}$, implies ${\sf AD}^{L({\mathbb R})}$ and much more, see \cite{john}, \cite{stacking}, and \cite[Chapter 12]{trang-sargsyan}.
This paper produces a proof of the following result.

\begin{theorem}\label{main_thm}
Assume {\em Martin's Maximum}${}^{++}$. Then Woodin's ${\mathbb P}_{\rm max}$-axiom $(*)$ holds true.
\end{theorem}

Our key new idea is ($\Sigma$.8) on page \pageref{(C.9)} below.


P.\ Larson, see \cite{plarson} and \cite{plarson2}, has shown that ${\sf MM}^{+\omega}$ is consistent with $\lnot (*)$ 
relative to a supercompact limit of supercompact
cardinals.

Throughout our entire paper, ``$\omega_1$'' 
will {\em always} denote $\omega_1^V$, the $\omega_1$ of $V$.  

Let us fix throughout this paper some $A \subset \omega_1$ such that $\omega_1^{L[A]} = \omega_1$. Let us 
define $g_A$ as the set of all ${\mathbb P}_{\rm max}$ conditions $p=(N;\in,I,a)$
such that there is a generic iteration $$(N_i,\sigma_{ij} \colon i \leq j \leq \omega_1)$$
of $p = N_0$ of length $\omega_1+1$ such that if we write $N_{\omega_1} = (N_{\omega_1};\in,I^*, a^*)$,\footnote{Here and elsewhere we often confuse a model with its underlying universe.} then $I^* = ({\sf NS}_{\omega_1})^V \cap N_{\omega_1}$ and $a^* = A$.

\begin{lemma}\label{folklore} \textbf{\em (Woodin)} Assume that ${\sf NS}_{\omega_1}$ is saturated and that ${\cal P}(\omega_1)^\#$ exists.
\begin{enumerate}
\item[(1)] $g_A$ is a filter.
\item[(2)] If $g_A$ is ${\mathbb P}_{\rm max}$-generic over $L({\mathbb R})$,
then ${\cal P}({\omega_1}) \subset L({\mathbb R})[g]$.
\end{enumerate}
\end{lemma}

{\sc Proof.} This routinely follows from the proof of \cite[Lemma 3.12 and Corollary 3.13]{hugh} and from \cite[Lemma 3.10]{hugh}. \hfill $\square$

\bigskip
One may also use {\sf BMM} plus ``${\sf NS}_{\omega_1}$ is precipitous'' to show that
$g_A$ is a filter, this is by the proof from \cite{cs}.

Let $\Gamma \subset \bigcup_{k<\omega} \, {\cal P}({\mathbb R}^k)$. We say that $\Gamma$ is {\em productive} iff for all $k<\omega$ and all $D \in \Gamma \cap {\cal P}({\mathbb R}^{k+2})$, if $D$ is universally Baire (see \cite{FMW}) as being witnessed by the trees $T$ and $U$ on ${}^{k+2} \omega \times {\rm OR}$, i.e., $D = p[T]$ and for all posets ${\mathbb P}$,
\begin{eqnarray}
\Vdash_{\mathbb P} \, p[U] = {\mathbb R}^{k+2} \setminus p[T]{\rm , }
\end{eqnarray}
and if 
\begin{eqnarray}\label{U-tilde}
{\tilde U} = \{ (s \upharpoonright (k+1), (s(k+1),t)) \colon (s,t) \in U \}{\rm , }
\end{eqnarray}
so that $(x_0, \ldots , x_{k}) \in p[{\tilde U}]$ iff there is some $y$ such that
$(x_0, \ldots , x_{k},y) \in p[U]$,
then 
there is a tree ${\tilde T}$ on ${}^{k+1} \omega \times {\rm OR}$ such that for all posets ${\mathbb P}$, 
\begin{eqnarray}
\Vdash_{\mathbb P} \, p[{\tilde U}] = {\mathbb R}^{k+1} \setminus p[{\tilde T}].
\end{eqnarray}

Let us denote by $\Gamma^\infty$ the collection of all $D \in \bigcup_{k<\omega} \, {\cal P}({\mathbb R}^k)$ which are universally Baire. If $D \in \Gamma^\infty$, then there is an 
unambiguous version of $D$ in any forcing extension of $V$, which as 
usual we 
denote by $D^*$. (\ref{U-tilde}) then means that 
if $D=p[U]$ and $E = p[{\tilde U}]$, then in any forcing extension of $V$,
$E^* = \exists^{\mathbb R} \, D^*$.

If $\Gamma \subset \Gamma^\infty$ is productive and if $D \in \Gamma$, then any projective statement about $D$ is absolute between $V$ and any forcing extension of $V$,\footnote{This seems to be wrong if we just assume $\Gamma \subset \Gamma^\infty$, but the hypothesis that $\Gamma$ be productive is crossed out.} i.e., if $\varphi$ is projective, $x_1$, $\ldots$, $x_k \in {\mathbb R}$, and ${\mathbb P}$ is any poset, then
$$V \models \varphi(x_1, \ldots , x_k, D) \ \Longleftrightarrow \ \Vdash_{\mathbb P} \, \varphi({\check x}_1, \ldots , {\check x}_k, {D}^*).$$

By a theorem of Woodin,
see e.g.\ \cite[Theorem 1.2]{forcing-free}, combined with the key result of Martin and Steel
in \cite{proj-determinacy}, the pointclass $\Gamma^\infty$ is productive under the
hypothesis that there is a proper class of Woodin cardinals. 

The proof from \cite{john} produces the result that under {\sf PFA}, the universe is closed under the operation $X \mapsto M_\omega^\#(X)$, which implies that every set of reals in $L({\mathbb R})$ is universally Baire and that $\bigcup_{k<\omega} \, {\cal P}({\mathbb R}^k) \cap L({\mathbb R})$ is productive. See e.g.\ \cite[Section 3, pp.\ 187f.]{both} on the relevant argument.
Therefore, in the light of Lemma \ref{folklore},
Theorem \ref{main_thm} 
follows from the following more general statement.

\begin{theorem}\label{main_thm_gen}
Let $\Gamma \subset \bigcup_{k<\omega} \, {\cal P}({\mathbb R}^k)$. Assume that
\begin{enumerate}
\item[(i)] $\Gamma = \bigcup_{k<\omega} \, {\cal P}({\mathbb R}^k) \cap L(\Gamma,{\mathbb R})$,
\item[(ii)] $\Gamma \subset \Gamma^\infty$,
\item[(iii)] $\Gamma$ is productive, and
\item[(iv)] Martin's Maximum${}^{++}$ holds true.
\end{enumerate}
Then $g_A$ is ${\mathbb P}_{\rm max}$-generic over $L(\Gamma,{\mathbb R})$.\footnote{In the presence of a proper class of Woodin cardinals, hypotheses (i), (ii), and (iv) then give $(*)_\Gamma$, see \cite[Definition 4.1]{both}.}
\end{theorem}

In the light of Lemma \ref{folklore} and \cite[Corollary 17]{FMS}, Theorem
\ref{main_thm_gen} readily follows from the following via a standard application
of ${\sf MM}^{++}$.

\begin{lemma}\label{main_claim}
Let $\Gamma \subset \bigcup_{k<\omega} \, {\cal P}({\mathbb R}^k)$. Assume that
\begin{enumerate}
\item[(i)] $\Gamma = \bigcup_{k<\omega} \, {\cal P}({\mathbb R}^k) \cap L(\Gamma,{\mathbb R})$,
\item[(ii)] $\Gamma \subset \Gamma^\infty$,
\item[(iii)] $\Gamma$ is productive, and
\item[(iv)] ${\sf NS}_{\omega_1}$ is saturated.\footnote{We could weaken this hypothesis to ``${\sf NS}_{\omega_1}$ is precipitous.''}
\end{enumerate}

Let $D \subset {\mathbb P}_{\rm max}$ be open dense, $D \in L(\Gamma,{\mathbb R})$.\footnote{By hypothesis, $D$ is then universally Baire in the codes, so that
there is an 
unambiguous version of $D$ in any forcing extension of $V$, which again 
we 
denote by $D^*$.}
There is then a stationary set preserving forcing ${\mathbb P}$ such that in $V^{\mathbb P}$
there is some $p = (N;\in,I^*,a^*) \in D^*$ and some generic iteration $$(N_i,\sigma_{ij} \colon i \leq j \leq \omega_1)$$
of $p = N_0$ of length $\omega_1+1$ such that if we write $N_{\omega_1} = (N_{\omega_1};\in,I^*, a^*)$, then $I^* = ({\sf NS}_{\omega_1})^{V^{\mathbb P}} \cap N_{\omega_1}$ and $a^* = A$.
\end{lemma}

The attentive reader will notice that we don't need the full power of ${\sf MM}^{++}$ 
in order to derive Theorem \ref{main_thm_gen} from Lemma \ref{main_claim}, 
the hypothesis that
$D$-${\sf BMM}^{++}$ holds true for all $D \in \Gamma^\infty$ would suffice, see
\cite[Definition 10.123]{hugh}.  By the proof of \cite[Theorem 2.7]{david-ralf}, $(*)$ is then actually {\em equivalent} to a version of {\sf BMM}; we state this as follows. 

\begin{theorem}
Let $\Gamma \subset \bigcup_{k<\omega} \, {\cal P}({\mathbb R}^k)$. Assume that
\begin{enumerate}
\item[(i)] $\Gamma = \bigcup_{k<\omega} \, {\cal P}({\mathbb R}^k) \cap L(\Gamma,{\mathbb R})$,
\item[(ii)] $\Gamma \subset \Gamma^\infty$,
\item[(iii)] $\Gamma$ is productive, and
\item[(iv)] ${\sf NS}_{\omega_1}$ is saturated.\footnote{Again, we could weaken this hypothesis to ``${\sf NS}_{\omega_1}$ is precipitous.''}
\end{enumerate}
The following statements are then equivalent.
\begin{enumerate}
\item[(1)] $D$-${\sf BMM}^{++}$ holds true for all $D \in \Gamma$. 
\item[(2)] $g_A$ is ${\mathbb P}_{\rm max}$-generic over $L(\Gamma,{\mathbb R})$.
\end{enumerate}
\end{theorem}
\begin{theorem} Assume that there is a proper class of Woodin cardinals. 
The following statements are then equivalent.
\begin{enumerate}
\item[(1)] $D$-${\sf BMM}^{++}$ holds true for all 
$D \in {\cal P}({\mathbb R}) \cap L({\mathbb R})$.
\item[(2)] $(*)$.
\end{enumerate}
\end{theorem}

Our next section is entirely devoted to a proof of Lemma \ref{main_claim}.

The authors thank Andreas Lietz for his comments on earlier versions of this paper.

\section{The forcing.}

Let us assume throughout the hypotheses of Lemma \ref{main_claim}.
We aim to verify its conclusion.

Let us fix $D \subset {\mathbb P}_{\rm max}$, an open dense set in $L(\Gamma,{\mathbb R})$.
By hypotheses (ii) and (iii) in the statement of Lemma \ref{main_claim} we will have that
\begin{enumerate}
\item[(D.1)] $V^{{\rm Col}(\omega,\omega_2)} \models$ ``$D^*$ is an open dense
subset of ${\mathbb P}_{\rm max}$.''
\end{enumerate}
Let us identify $D$ with a canonical set of reals coding the elements of $D$,\footnote{We will have to spell out a bit more precisely below in which way we aim to have the elements of $p[T]$ code the elements of $D$, see ($\Sigma$.5) below.} and
let $T \in V$ be a
tree on $\omega \times 2^{{\aleph_2}}$ such that 
\begin{enumerate}
\item[(D.2)] $V^{{\rm Col}(\omega,\omega_2)} \models D^*=p[T]$. 
\end{enumerate}

Let us write 
\begin{eqnarray} \kappa = (2^{\aleph_2})^+ {\rm , }
\end{eqnarray}
so that $T \in H_\kappa$.
Let $d$ be ${\rm Col}(\kappa,\kappa)$-generic over $V$. In $V[d]$, let
$({\bar A}_\lambda \colon \lambda<\kappa)$ be a $\Diamond_\kappa$-sequence, i.e., 
for all ${\bar A} \subset \kappa$, $\{ \lambda<\kappa \colon {\bar A} \cap \lambda = {\bar A}_\lambda \}$ is stationary.
Also, let $c \colon \kappa \rightarrow H_\kappa^{V} = H_\kappa^{V[d]}$,
$c \in V[d]$, be bijective.
For $\lambda < \kappa$, let 
\begin{eqnarray}
Q_\lambda = c \mbox{''} \lambda \mbox{ and } A_\lambda = c \mbox{''} {\bar A}_\lambda.
\end{eqnarray} 
Let $C \subset \kappa$ be club such that for all $\lambda \in C$,
\begin{enumerate}
\item[(i)] $Q_\lambda$
is transitive, 
\item[(ii)] $\{ T , ((H_{\omega_2})^V ; \in , ({\sf NS}_{\omega_1})^V ,A) \} \cup
2^{\aleph_2} \subset Q_\lambda$,
\item[(iii)] $Q_\lambda \cap {\rm OR} = \lambda$ (so that $c \upharpoonright \lambda \colon
\lambda \rightarrow Q_\lambda$ is bijective), and 
\item[(iv)] $({\cal Q}_\lambda;\in) \prec
(H_\kappa;\in)$.
\end{enumerate}
In $V[d]$, for all $P$, $B \subset H_\kappa$, the set of all $\lambda \in C$ such that
$$(Q_\lambda;\in,P \cap Q_\lambda , B \cap Q_\lambda) \prec (H_\kappa;\in,P,B)$$
is club, and the set of all $\lambda \in C$ such that $B \cap Q_\lambda = A_\lambda$ is stationary, so that 
\begin{enumerate}
\item[($\Diamond$)] For all $P$, $B \subset H_\kappa$ the set
$$\{ \lambda \in C \colon (Q_\lambda;\in , P \cap Q_\lambda , A_\lambda) \prec (H_\kappa;\in , P , B) \}$$
is stationary.
\end{enumerate}

We shall sometimes also write $Q_\kappa=H_\kappa$.

We shall now go ahead and 
produce a stationary set preserving forcing ${\mathbb P} \in V[d]$ which adds some
$p \in D^*$ and some generic iteration $$(N_i,\sigma_{ij} \colon i \leq j \leq \omega_1)$$
of $p = N_0$ such that if we write $N_{\omega_1} = (N_{\omega_1};\in,I^*, a^*)$, then $I^* = ({\sf NS}_{\omega_1})^{V[d]^{\mathbb P}} \cap N_{\omega_1}$ and $a^* = A$. As the forcing ${\rm Col}(\kappa,\kappa)$ which added $d$ is certainly stationary set preserving, this will verify Lemma \ref{main_claim}.

${\sf NS}_{\omega_1}$ is still saturated in $V[d]$ and (D.1) and (D.2) are still true in $V[d]$, so that in order to simplify our notation, we shall in what follows confuse $V[d]$ with $V$, i.e., pretend that in addition to ``${\sf NS}_{\omega_1}$ is saturated'' 
plus (D.1) and (D.2), ($\Diamond$) is also true in $V$. 

Working under these hypotheses, we shall now
recursively define a $\subset$-increasing and continuous chain of forcings ${\mathbb P}_\lambda$ for all 
$\lambda \in C \cup \{ \kappa \}$. The forcing ${\mathbb P}$ will be ${\mathbb P}_\kappa$.

Assume that $\lambda \in C \cup \{ \kappa \}$ and ${\mathbb P}_\mu$ has already 
been defined in such a way that ${\mathbb P}_\mu \subset Q_\mu$
for all $\mu \in C \cap \lambda$.

We shall be interested in objects ${\mathfrak C}$ which exist in some outer model\footnote{$W$ is an outer model iff $W$ is a transitive model of {\sf ZFC} with
$W \supset V$ and which has the same ordinals as $V$; in other words, $W$ is an 
outer model iff $V$ is an inner model of $W$.}
and which have the following properties.
\begin{eqnarray}\label{certificate}
{\mathfrak C} = \langle M_i , \pi_{ij} , N_i , \sigma_{ij} \colon i \leq j \leq \omega_1 \rangle , 
\langle (k_n , \alpha_n ) \colon n<\omega \rangle ,
\langle \lambda_\delta , X_\delta \colon \delta \in K \rangle {\rm , }
\end{eqnarray}
and
\begin{enumerate}
\item[(C.1)] $M_0$, $N_0 \in {\mathbb P}_{\rm max}$,
\item[(C.2)] $x = \langle k_n \colon n<\omega \rangle$ is a real code for $N_0$ and
$\langle (k_n , \alpha_n ) \colon n<\omega \rangle \in [T]$,
\item[(C.3)] $\langle M_i , \pi_{ij} \colon i \leq j \leq \omega_1^{N_0} \rangle \in N_0$ is a generic iteration of $M_0$ which witnesses that $N_0
< M_0$ in ${\mathbb P}_{\rm max}$,
\item[(C.4)] $\langle N_i , \sigma_{ij} \colon i \leq j \leq \omega_1 \rangle$ is a generic iteration of $N_0$ such that if $$N_{\omega_1} = (N_{\omega_1};\in,I^*,A^*){\rm , }$$ then 
$A^* = A$,\footnote{There is no requirement on $I^*$ matching the non-stationary ideal of some model in which ${\mathfrak C}$ exists.\label{no-requirement}}
\item[(C.5)] $\langle M_i , \pi_{ij} \colon i \leq j \leq \omega_1 \rangle = \sigma_{0 \omega_1}( \langle M_i , \pi_{ij} \colon i \leq j \leq \omega_1^{N_0}\rangle )$ and \begin{eqnarray}\label{Momega1}
M_{\omega_1} = ((H_{\omega_2})^V ; \in , ({\sf NS}_{\omega_1})^V ,A){\rm , }
\end{eqnarray}
\item[(C.6)] $K \subset {\omega_1}$, 
\end{enumerate}
and for all $\delta \in K$,
\begin{enumerate}
\item[(C.7)] $\lambda_\delta<\lambda$, and if $\gamma<\delta$ is in $K$, then 
$\lambda_\gamma<\lambda_\delta$ and 
$X_\gamma \cup \{ \lambda_\gamma \} \subset X_\delta$, and
\item[(C.8)] $X_\delta \prec (Q_{\lambda_\delta};\in,{\mathbb P}_{\lambda_\delta},A_{\lambda_\delta})$ and $X_\delta \cap \omega_1 = \delta$.
\end{enumerate}

We need to define a language ${\cal L}$ (independently from $\lambda$)
whose formulae will be able to describe ${\mathfrak C}$ with the above properties by producing the 
models $M_i$ and $N_i$, $i <\omega_1$, as term models out
of equivalence classes of terms of the form ${\dot n}$, $n<\omega$.
The language ${\cal L}$ will have the the following constants.
\begin{align*} 
{\dot T} & \mbox{ \ \ \ intended to denote } T \\
x \mbox{ for every } x \in H_\kappa  & \mbox{ \ \ \ intended to denote } x \mbox{ itself } \\ 
{\dot n} \mbox{ for every } n<\omega & \mbox{ \ \ \ as terms for elements of } M_i \mbox{ and } N_i, i < \omega_1 \\
{\dot M}_i \mbox{ for } i<\omega_1 & \mbox{ \ \ \ intended to denote } M_i \\
{\dot \pi}_{ij} \mbox{ for } i \leq j \leq \omega_1 & \mbox{ \ \ \ intended to denote } \pi_{ij} \\
{\dot {\vec M}} & \mbox{ \ \ \ intended to denote } (M_j , \pi_{jj'} \colon j \leq j' \leq 
\omega_1^{N_i}) \mbox{ for } i < \omega_1 \\
{\dot N}_i \mbox{ for } i < \omega_1 & \mbox{ \ \ \ intended to denote } N_i \\
{\dot \sigma}_{ij} \mbox{ for } i \leq j < \omega_1 & \mbox{ \ \ \ intended to denote } \sigma_{ij} \\
{\dot a} & \mbox{ \ \ \ intended to denote the distinguished $a$-predicate of } M_i, N_i, i<\omega_1 \\
{\dot I} 
& \mbox{ \ \ \ intended to denote the distinguished ideal of } N_i, i<\omega_1\\
{\dot X}_\delta \mbox{ for } \delta<\omega_1 & \mbox{ \ \ \ intended to denote } X_\delta.
\end{align*}
Formulae of ${\cal L}$ will be of the following form.
\begin{gather*} 
\ulcorner {\dot N}_i \models \varphi(\xi_1, \ldots , \xi_k , {\dot n}_1 , \ldots , {\dot n}_\ell, {\dot a}, {\dot I} , 
{\dot M}_{j_1}, \ldots, {\dot M}_{j_m} , {\dot \pi}_{q_1 r_1} , \ldots , {\dot \pi}_{q_s r_s} , {\dot {\vec M}}) \urcorner \\
\mbox{for } i<\omega_1,
\xi_1, \ldots , \xi_k < \omega_1, n_1, \ldots , n_\ell < \omega, j_1, \ldots, j_m < \omega_1,
q_1 \leq r_1 < \omega_1, \ldots, q_s \leq r_s < \omega_1 \\
\ulcorner {\dot \pi}_{i \omega_1}({\dot n})=x \urcorner  \mbox{ \ \ \ for }
i < \omega_1 \mbox{ and } x \in H_{\omega_2} \\
\ulcorner {\dot \pi}_{\omega_1 \omega_1}(x)=x \urcorner  \mbox{ \ \ \ for }
x \in H_{\omega_2} \\ 
\ulcorner {\dot \sigma}_{ij}({\dot n})={\dot m} \urcorner  \mbox{ \ \ \ for }
i \leq j < \omega_1, n, m<\omega \\
\ulcorner ({\vec u},{\vec \alpha}) \in {\dot T} \urcorner  \mbox{ \ \ \ 
for } {\vec u} \in {}^{<\omega} \omega \mbox{ and } {\vec \alpha} 
\in {}^{<\omega} (2^{\aleph_2}) \\
\ulcorner \delta \mapsto {\bar \lambda} \urcorner  \mbox{ \ \ \ for } \delta<\omega_1,
{\bar \lambda}<\kappa \\
\ulcorner x \in {\dot X}_\delta \urcorner  \mbox{ \ \ \ for } \delta<\omega_1, x \in H_\kappa
\end{gather*}

Let us write ${\cal L}^\lambda$ for the collection of all ${\cal L}$-formulae except for the 
formulae which mention elements outside of $Q_\lambda$, i.e., except for the formulae
of the form $\ulcorner \delta \mapsto {\bar \lambda} \urcorner$
for $\delta < \omega_1$ and ${\lambda} \leq {\bar \lambda} < \kappa$ and $\ulcorner x \in {\dot X}_\delta \urcorner$ for $\delta < \omega_1$ and $x
\in H_\kappa \setminus Q_\lambda$.
We may and shall assume that ${\cal L}$ is built in a canonical way so that 
${\cal L}^\lambda \subset Q_\lambda$.

We say that the objects ${\mathfrak C}$ as in
(\ref{certificate}) are {\em pre-certified} by a collection $\Sigma$ of ${\cal L}^\lambda$-formulae 
if and only if (C.1) through (C.8) are satisfied by ${\mathfrak C}$
and there are surjections $e_i \colon \omega \rightarrow
N_i$ 
for $i<\omega_1$
such that the following hold true.
\begin{enumerate}
\item[($\Sigma$.1)] 
$\ulcorner {\dot N}_i \models \varphi(\xi_1, \ldots , \xi_k , {\dot n}_1 , \ldots , {\dot n}_\ell, {\dot a}, {\dot I} , 
{\dot M}_{j_1}, \ldots, {\dot M}_{j_m} , {\dot \pi}_{q_1 r_1} , \ldots , {\dot \pi}_{q_s r_s} , {\dot {\vec M}})
\urcorner \in \Sigma$ iff \newline $i<\omega_1$,
$\xi_1$, $\ldots$, $\xi_k \leq \omega_1^{N_i}$, $n_1$, $\ldots$ , $n_\ell < \omega$, $j_1$, $\ldots$, $j_m \leq \omega_1^{N_i}$,
$q_1 \leq r_1 \leq \omega_1^{N_i}$, $\ldots$, $q_s \leq r_s \leq \omega_1^{N_i}$ and \newline
$N_i \models \varphi(\xi_1, \ldots , \xi_k , e_i(n_1) , \ldots , e_i(n_\ell), A \cap \omega_1^{N_i}, I^{N_i} , M_{j_1}, \ldots
, M_{j_m}, \pi_{q_1 r_1}, \ldots , \pi_{q_s r_s}, {\vec M})$, where $I^{N_i}$ is the distinguished ideal of $N_i$ and ${\vec M}=\langle M_j , \pi_{jj'} \colon
j \leq j' \leq \omega_1^{N_i})$,
%
%
\item[($\Sigma$.2)] $\ulcorner {\dot \pi}_{i \omega_1}({\dot n})=x \urcorner \in \Sigma$ iff $i < \omega_1$, $n<\omega$, and $\pi_{i \omega_1}(e_i(n))=x$,
\item[($\Sigma$.3)] $\ulcorner {\dot \pi}_{\omega_1 \omega_1}(x)=x \urcorner \in \Sigma$ iff $x \in H_{\omega_2}$,
\item[($\Sigma$.4)] $\ulcorner {\dot \sigma}_{ij}({\dot n})={\dot m} \urcorner \in \Sigma$ iff $i \leq j < \omega_1$, $n$, $m<\omega$, and $\sigma_{ij}(e_i(n))=e_j(m)$,
\item[($\Sigma$.5)] letting $F$ with ${\rm dom}(F)=\omega$ be the monotone enumeration 
of the G\"odel numbers of all $\ulcorner {\dot N}_0 \models \varphi({\dot n}_1, \ldots , {\dot n}_\ell, {\dot a}, {\dot I})
\urcorner$ with $\ulcorner {\dot N}_0 \models \varphi({\dot n}_1, \ldots , {\dot n}_\ell, {\dot a}, {\dot I})
\urcorner \in \Sigma$, we have that $\ulcorner ({\vec u},{\vec \alpha}) \in {\dot T} \urcorner \in \Sigma$ iff
there is some $n<\omega$ such that $\langle {\vec u},{\vec \alpha} \rangle 
= \langle (F(m),\alpha_m ) \colon m< n \rangle$ and $F(m)=k_m$ for all $m<n$, 
\item[($\Sigma$.6)] $\ulcorner \delta \mapsto {\bar \lambda} \urcorner \in \Sigma$ iff
$\delta \in K$ and ${\bar \lambda} = \lambda_\delta$, and
\item[($\Sigma$.7)] $\ulcorner {x} \in {\dot X}_\delta \urcorner \in \Sigma$ iff
$\delta \in K$ and $x \in X_\delta$.
\end{enumerate}

We say that the objects ${\mathfrak C}$ as in
(\ref{certificate}) are {\em certified} by a collection $\Sigma$ of formulae 
if and only if ${\mathfrak C}$ is pre-certified by $\Sigma$ and in addition,
\begin{enumerate}
\item[($\Sigma$.8)]\label{(C.9)} if $\delta \in K$, then 
$[\Sigma]^{<\omega} \cap X_\delta \cap
E \not= \emptyset$ for every $E \subset {\mathbb P}_{\lambda_\delta}$ which is dense in ${\mathbb P}_{\lambda_\delta}$ and definable over the structure $$(Q_{\lambda_\delta}; \in , {\mathbb P}_{\lambda_\delta}  , A_{\lambda_\delta} )$$ from parameters in $X_\delta$.\footnote{Equivalently,
$[\Sigma]^{<\omega} \cap
E \not= \emptyset$ for every $E \subset {\mathbb P}_{\lambda_\delta} \cap 
X_\delta$ which is dense in ${\mathbb P}_{\lambda_\delta} \cap 
X_\delta$ and definable over the structure $$(X_\delta; \in , {\mathbb P}_{\lambda_\delta} \cap 
X_\delta , A_{\lambda_\delta} \cap X_\delta )$$ from parameters in $X_\delta$.}
\end{enumerate}

By way of definition, we call ${\mathfrak C}$ as in (\ref{certificate}) a {\em semantic
certificate} iff 
there is a 
collection $\Sigma$ of formulae such that ${\mathfrak C}$ is certified by $\Sigma$.
We call $\Sigma$ a {\em syntactic certificate}
iff there is a semantic certificate ${\mathfrak C}$ such that ${\mathfrak C}$ is certified by
$\Sigma$. Given a syntactic certificate $\Sigma$, there is a {\em unique} semantic
certificate ${\mathfrak C}$ such that ${\mathfrak C}$ is certified by $\Sigma$.
Even though it is obvious how to construct ${\mathfrak C}$ from $\Sigma$, in the proof of Lemma \ref{generics_give_certificates} below we \label{details} will provide details on how to derive a semantic certificate from a given $\Sigma$.

Let $\Sigma \cup p$ be a set of formulae, where $p$ is finite.
We say that $p$
is {\em certified by} 
$\Sigma$ if and only if
there is some (unique) ${\mathfrak C}$ as in (\ref{certificate})
such that 
${\mathfrak C}$ is certified by 
$\Sigma$ and   
\begin{enumerate}
\item[($\Sigma$.9)] $p \in [\Sigma]^{<\omega}$.
\end{enumerate}
We may also say that $p$ is {\em certified by} ${\mathfrak C}$ as in (\ref{certificate}) iff there is some $\Sigma$ such that ${\mathfrak C}$ and $p$ are both certified by $\Sigma$ -- and we will then also refer to $\Sigma$ as a syntactical certificate for $p$ and to ${\mathfrak C}$ as the associated semantical certificate.

We are then ready to define the forcing ${\mathbb P}_\lambda$. 
We say that $p \in {\mathbb P}_\lambda$ if and only if 
\begin{eqnarray}\label{in_gen_ext}
V^{{\rm Col}(\omega,\lambda)} \models \mbox{``There is a set $\Sigma$ of ${\cal L}^\lambda$-formulae such that } p \mbox{ is certified by } \Sigma \mbox{.''}
\end{eqnarray}

Let $p$ be a finite set of formulae of ${\cal L}^\lambda$. By the homogeneity of ${\rm Col}(\omega,\lambda)$, if there is some $h$ which is 
${\rm Col}(\omega,\lambda)$-generic over $V$ and there is some 
$\Sigma
\in V[h]$
such that $p$ is certified by $\Sigma$, then for all $h$ which are ${\rm Col}(\omega,\lambda)$-generic over $V$ there is some $\Sigma \in V[h]$ such that $p$ is certified by $\Sigma$.
It is then easy to see that $\langle {\mathbb P}_\lambda \colon \lambda \in C \cup
\{ \kappa \} \rangle$ is definable over $V$ from $\langle A_\lambda \colon
\lambda < \kappa \rangle$
and $C$, and is hence an element of $V$.

Again let $p$ be a finite set of formulae of ${\cal L}^\lambda$. By $\Sigma^1_1$ absoluteness, if there is any outer model in which there is some $\Sigma$ which
certifies $p$, then there is some $\Sigma \in V^{{\rm Col}(\omega,\lambda)}$
which certifies $p$.\footnote{In fact, if $P$ is a transitive model of {\sf KP} plus the axiom
{\em Beta} with $(Q_\lambda; \langle A_{\bar \lambda} \colon {\bar \lambda} < \lambda \rangle ) \in P$ and if $p \in {\mathbb P}_\lambda$,
then there is some $\Sigma \in P^{{\rm Col}(\omega,\lambda)}$ which certifies $p$.}  This simple observation is important in the verification
that ${\mathbb P}_\lambda$ is actually non-empty, cf.\ Lemma \ref{P_is_nonempty}, 
and in the proof of Lemma
\ref{P_stat_pres}.

It is easy to see that \begin{enumerate}
\item[(i)] ${\mathbb P}={\mathbb P}_\kappa \subset H_\kappa$,
\item[(ii)] if ${\bar \lambda} < \lambda$ are both in $C \cup \{ \kappa \}$, then ${\mathbb P}_{\bar \lambda} \subset {\mathbb P}_\lambda$, and
\item[(iii)] if $\lambda \in C \cup \{ \kappa \}$ is a limit point of 
$C \cup \{ \kappa \}$, then ${\mathbb P}_\lambda = \bigcup_{{\bar \lambda} \in C \cap \lambda} \, {\mathbb P}_{\bar \lambda}$,
\end{enumerate}
so that there is some club $D \subset C$ such that for all $\lambda \in D$,
$${\mathbb P}_\lambda = {\mathbb P} \cap Q_\lambda.$$ Hence ($\Diamond$) gives us the following.
\begin{enumerate}\label{diamond-thing}
\item[($\Diamond({\mathbb P})$)] For all $B \subset H_\kappa$ the set
$$\{ \lambda \in C \colon (Q_\lambda;\in , {\mathbb P}_\lambda , A_\lambda) \prec (H_\kappa;\in , {\mathbb P} , B) \}$$
is stationary.
\end{enumerate}

The first one of the following lemmas is entirely trivial.

\begin{lemma}\label{+++}
Let $\Sigma$ be a syntactic certificate, and let $p$, $q \in [\Sigma]^{<\omega}$. Then $p$ and $q$ are compatible conditions in ${\mathbb P}$. 
\end{lemma}

\begin{lemma}\label{P_is_nonempty}
Let $\lambda \in C \cup \{ \kappa \}$. Then $\emptyset \in {\mathbb P}_\lambda$.
\end{lemma}

{\sc Proof.} See the proof of \cite[Theorem 2.8]{david-ralf}, or the proof of
\cite[Theorem 4.2]{both}. Notice that for all $\lambda \in C \cup \{ \kappa \}$,
$\emptyset \in {\mathbb P}_\lambda$ iff $\emptyset \in {\mathbb P}$.

Let
$h$ be ${\rm Col}(\omega,\omega_2)$-generic over $V$, and write $\rho=\omega_3^V=
\omega_1^{V[h]}$. Inside $V[h]$, $$((H_{\omega_2})^V;\in ({\sf NS}_{\omega_1})^V , A)$$ is a ${\mathbb P}_{\rm max}$ condition, call it $p$.
Let $q \in ({\mathbb P}_{\rm max})^{V[h]}$, $q < p$, $q \in D^*$, cf.\ (D.1).
Let $q = N_0 = (N_0; \in , I , a)$. 
Let $(M_i, \pi_{ij} \colon i \leq j \leq \omega_1^{N_0}) \in N_0$ be the unique generic iteration of $p$ which witnesses $q < p$. 

Let $(N_i,\sigma_{ij} \colon i \leq j \leq \rho) \in V[h]$ be a generic iteration of
$N_0$ such that $\rho=\omega_1^{N_\rho}$.\footnote{If we wished, then we could even arrange that
writing $N_\rho = (N_\rho;\in,I^*,a^*)$, we have that $I^* = ({\sf NS}_\rho)^{V[h]}
\cap N_\rho$, but this is not relevant here; cf.\ footnote \ref{no-requirement}.} Let 
\begin{eqnarray}\label{it_of_p}
(M_i , \pi_{ij} \colon i \leq j \leq \rho)=
\sigma_{0 \rho}((M_i, \pi_{ij} \colon i \leq j \leq \omega_1^{N_0}))
\end{eqnarray}
We may lift (\ref{it_of_p}) to a generic iteration 
\begin{eqnarray}\label{it_of_p}
(M_i^+ , \pi^+_{ij} \colon i \leq j \leq \rho)
\end{eqnarray}
of $V$. Let us write $M=M^+_\rho$ and $\pi = \pi^+_{0 \rho}$.

Let $\langle k_n , \alpha_n \colon n<\omega \rangle$ be such that $x
= \langle k_n \colon n<\omega \rangle$ is
a real code for $N_0$ \`a la ($\Sigma$.5) and $\langle (k_n , \alpha_n) \colon n<\omega \rangle \in [T]$.
We then clearly have that $\langle (k_n , \pi(\alpha_n)) \colon n<\omega \rangle \in [\pi(T)]$.

It is now easy to see that
\begin{eqnarray}\label{certificate_at_rho}
{\mathfrak C} = \langle M_i , \pi_{ij} , N_i , \sigma_{ij} \colon i \leq j \leq \rho \rangle , 
\langle (k_n , \pi(\alpha_n) ) \colon n<\omega \rangle ,
\langle \rangle
\end{eqnarray} 
certifies $\emptyset$, construed as the empty set of $\pi({\cal L}^\kappa)$ formulae: as the third
component $\langle \rangle$ of ${\mathfrak C}$ in (\ref{certificate_at_rho}) is empty, {\em any} set of surjections $e_i \colon \omega \rightarrow N_i$, $i<\omega_1$, will induce a syntactic certificate for $\emptyset$ whose associated sematic certificate is ${\mathfrak C}$. By $\Sigma^1_1$ absoluteness, there is then some
${\mathfrak C} \in M^{{\rm Col}(\omega,\pi(\omega_2))}$ as in (\ref{certificate_at_rho}) which certifies $\emptyset$ so that 
$\emptyset \in \pi({\mathbb P})$. By the elementarity of $\pi$,
then, there is some ${\mathfrak C} \in V^{{\rm Col}(\omega,\omega_2)}$
which certifies $\emptyset$, construed as the empty set of ${\cal L}^\kappa$ formulae. Hence 
$\emptyset \in {\mathbb P}$.
\hfill $\square$

\begin{lemma}\label{generics_give_certificates}
Let $\lambda \in C \cup \{ \kappa \}$.
Let $g \subset {\mathbb P}_\lambda$ be a filter such that $g \cap E \not= \emptyset$ for all dense $E \subset {\mathbb P}_\lambda$ which are definable over $(Q_\lambda;\in,{\mathbb P}_\lambda)$ from elements of $Q_\lambda$. Then 
$\bigcup g$
is a syntactic certificate.
\end{lemma}

{\sc Proof.}
It is obvious how to read off from $\bigcup g$ a candidate
$${\mathfrak C} = \langle M_i , \pi_{ij} , N_i , \sigma_{ij} \colon i \leq j \leq \omega_1 \rangle , 
\langle (k_n , \alpha_n ) \colon n<\omega \rangle ,
\langle \lambda_\delta , X_\delta \colon \delta \in K \rangle$$
like in (\ref{certificate}) for
a semantical certificate for $\bigcup g$. 
Let us be somewhat explicit, though. A variant of what is to come shows 
how to derive ${\mathfrak C}$ from a given syntactic certificate $\Sigma$,
where ${\mathfrak C}$ is unique such that $\Sigma$ certifies ${\mathfrak C}$, cf.\ the remark on p.\ \pageref{details}. 

For 
$i$, $j <\omega_1$ and 
$\tau$, $\sigma \in \{ {\dot n} \colon n<\omega \} \cup \omega_1$
define 
\begin{align*} 
\tau \sim_{i} \sigma & \mbox{ \ \ \ iff \ \ } \ulcorner {\dot N}_i \models \tau = \sigma \urcorner \in \bigcup g \\
(i,\tau) \sim_{\omega_1} (j,\sigma) & \mbox{ \ \ \ iff } i \leq j \wedge \exists
\rho \, \{ \ulcorner
{\dot \sigma}_{ij}(\tau)=\rho \urcorner , \ulcorner {\dot N}_j \models
\rho = \sigma \urcorner \} \subset \bigcup g\\
{} & \mbox{ \ \ \ \ or } j \leq i \wedge \exists
\rho \, \{ \ulcorner
{\dot \sigma}_{ji}(\sigma)=\rho \urcorner , \ulcorner {\dot N}_j \models
\rho = \tau \urcorner \} \subset \bigcup g\\
[\tau]_{i} & = \{ \sigma \colon \tau \sim_{i} \sigma \}  \\
[(i,\tau)] & = \{ (j,\sigma) \colon (i,\tau) \sim_{\omega_1} (j,\sigma) \}  \\
{M}_i & = \{ [\tau]_i \colon \tau \in \{ {\dot n} \colon n<\omega \} \cup \omega_1 \wedge \ulcorner {\dot N}_i \models \tau \in {\dot M}_i \urcorner \in \bigcup g \} \\
M_{\omega_1} & = (H_{\omega_2})^V \\
{N}_i & = \{ [\tau]_i \colon \tau \in \{ {\dot n} \colon n<\omega \} \cup \omega_1 \} \\
N_{\omega_1} & = \{ [i,\tau]_{\omega_1} \colon i<\omega_1 \wedge \tau \in {\dot M}_i \urcorner \in \bigcup g \} \\
[\tau]_{i} \, {\tilde \in}_i \, [\sigma]_{i} & \mbox{ \ \ \ iff \ \ } \ulcorner {\dot N}_i \models \tau \in \sigma \urcorner \in \bigcup g \\
[i,\tau] {\tilde \in}_{\omega_1} [j,\sigma] & \mbox{ \ \ \ iff }
i \leq j \wedge \exists
\rho \, \{ \ulcorner
{\dot \sigma}_{ij}(\tau)=\rho \urcorner , \ulcorner {\dot N}_j \models
\rho \in \sigma \urcorner \} \subset \bigcup g\\
{} & \mbox{ \ \ \ \ or } j \leq i \wedge \exists
\rho \, \{ \ulcorner
{\dot \sigma}_{ji}(\sigma)=\rho \urcorner , \ulcorner {\dot N}_j \models
\tau \in \rho \urcorner \} \subset \bigcup g\\
[\tau]_i \in I^{N_i} & \mbox{ \ \ \ iff \ \ } \ulcorner {\dot N}_i \models \tau \in {\dot I} \urcorner \in \bigcup g \\
[i,\tau] \in I^{N_{\omega_1}} & \mbox{ \ \ \ iff \ \ } [\tau]_i \in I^{N_i} \\
[\tau]_i \in a^{N_i} & \mbox{ \ \ \ iff \ \ } \ulcorner {\dot N}_i \models \tau \in {\dot a} \urcorner \in \bigcup g \\
[i,\tau] \in a^{N_{\omega_1}} & \mbox{ \ \ \ iff \ \ } [\tau]_i \in I^{N_i} \\ 
\pi_{ij}([\tau]_{i}) = [\sigma]_j & \mbox{ \ \ \ iff \ \ } \ulcorner {\dot N}_j \models 
{\dot \pi}_{ij}(\tau)=\sigma \urcorner \in \bigcup g \\
\pi_{i \omega_1}([\tau]_{i}) = x & \mbox{ \ \ \ iff \ \ } \ulcorner 
{\dot \pi}_{i \omega_1}(\tau)=x \urcorner \in \bigcup g \\
\pi_{\omega_1 \omega_1}(x) = x & \mbox{ \ \ \ iff \ \ } x \in (H_{\omega_2})^V \\
\sigma_{ij}([\tau]_{i}) = [\sigma]_j & \mbox{ \ \ \ iff \ \ } \ulcorner  
{\dot \sigma}_{ij}(\tau)=\sigma \urcorner \in \bigcup g \\
\sigma_{i \omega_1}([\tau]_i) = [i,\tau] & {} \\
%
(k,\alpha) = (k_n^g, \alpha_n^g) & \mbox{ \ \ \ iff } \exists {\vec u} \exists
{\vec \alpha} ( \ulcorner ({\vec u} , {\vec \alpha}) \in {\dot T} \urcorner
\in \bigcup g \wedge k = {\vec u}(n) \wedge  \alpha = {\vec \alpha}(n) )\\
\delta \in K^g & \mbox{ \ \ \ iff } \exists {\bar \lambda} \, \ulcorner \delta \mapsto {\bar \lambda} \urcorner \in \bigcup g\\
{\bar \lambda} = \lambda^g_\delta & \mbox{ \ \ \ iff } \delta \in K^g \wedge \ulcorner \delta \mapsto {\bar \lambda} \urcorner \in \bigcup g\\
x \in X_\delta^g & \mbox{ \ \ \ iff } \delta \in K^g \wedge \ulcorner x \in {\dot X}_\delta \urcorner \in \bigcup g
\end{align*}

We will first have that 
${\tilde \in}_0$ is wellfounded and that in fact (the transitive collapse of)
the structure $N_0 = (N_0;{\tilde \in}_0,a^{N_0},I^{N_0})$ is an iterable ${\mathbb P}_{\rm max}$ condition. This is true because  
straightforward density arguments give (C.2), i.e., that $\langle (k_n, \alpha_n) \colon n<\omega \rangle \in [T]$ and $\langle k_n \colon
n<\omega \rangle$ will code the theory of $N_0$ \`a la ($\Sigma$.5).

Another set of easy density arguments will give that $(N_i ,\sigma_{ij} \colon i \leq j \leq \omega_1)$
is a generic iteration of $N_0$, were we identify $N_i$ with the structure
$(N_i;{\tilde \in}_i,a^{N_i},I^{N_i})$. 
To verify this, let us first show:
\begin{claim}\label{claim2.4}
For each $i < \omega_1$ and for each $\xi \leq \omega_1^{N_i}$, $[\xi]_i$ represents 
$\xi$ in (the transitive collapse of the well-founded part of) the term model for $N_i$;
moreover, $a^{N_i} = A \cap \omega_1^{N_i}$. Hence $a^{N_{\omega_1}}=A$.
\end{claim}

{\sc Proof} of Claim \ref{claim2.4}. Straightforward arguments using ($\Sigma$.1) show that $[\xi]_i$ must always represent $\xi$ in $N_i$ as given by any certificate.
Claim \ref{claim2.4} then follows by straightforward density arguments. \hfill $\square$ (Claim \ref{claim2.4})

\bigskip
Similarily:

\begin{claim} Let $i<\omega_1$. 
$N_{i+1}$ is generated from ${\rm ran}(\sigma_{ii+1}) \cup \{ \omega_1^{N_i} \}$
in the sense that for every $x \in N_{i+1}$ there is some function $f \in {}^{\omega_1^{N_i}} 
(N_i) \cap N_i$ such that $x = \sigma_{ii+1}(f)(\omega_1^{N_i})$.
\end{claim}

\begin{claim}
Let $i<\omega_1$. $\{ X \in {\cal P}(\omega_1^{N_i}) \cap N_i \colon \omega_1^{N_i} \in 
\sigma_{ii+1}(X) \}$ is generic over $N_i$ for the forcing given by the $I^{N_i}$-positive sets.
\end{claim}

\begin{claim}
Let $i \leq \omega_1$ be a limit ordinal. For every $x \in N_i$ there is some $j<i$ and some ${\bar x} \in N_j$ such that $x = \sigma_{ji}({\bar x})$.
\end{claim}

$(N_i ,\sigma_{ij} \colon i \leq j \leq \omega_1)$
is then indeed a generic iteration of $N_0$. As $N_0$ is iterable, we may and shall identify $N_i$ with its transitive collapse, so that 
(C.4) holds true.

Another round of density arguments will show that
${\mathfrak C}$
satisfies (C.1), (C.3), (C.5), (C.6), and (C.7), where we identify $M_i$
with the structure $(M_i;\in,({\sf NS}_{\omega_1^{M_i}})^{M_i},A \cap \omega_1^{M_i})$.
Let us now verify (C.8) and (C.9).

As for (C.8), $X_\delta \cap \omega_1 = \delta$ for $\delta \in K$ is easy. Let $x_1$, $\ldots$, $x_k 
\in X_\delta$, $\delta \in K$.
Suppose that 
\begin{eqnarray}\label{in-the-model}
(Q_{\lambda_\delta};\in {\mathbb P}_{{\lambda_\delta}},
A_{{\lambda_\delta}}) \models \exists v \, \varphi(v,x_1,\ldots , x_k).
\end{eqnarray}
Let $p \in g$ be such that $\{ \ulcorner x_1 \in {\dot X}_\delta \urcorner , \ldots ,
\ulcorner x_k \in {\dot X}_\delta \urcorner , \ulcorner \delta \mapsto \lambda_\delta \urcorner
\} \subset p$. Let $q \leq p$, and let
$\Sigma$ be a syntactical certificate for $q$ whose associated semantical certificate is
$${\mathfrak C}' = \langle M'_i , \pi'_{ij} , N'_i , \sigma'_{ij} \colon i \leq j \leq \omega_1 \rangle , 
\langle (k'_n , \alpha'_n ) \colon n<\omega \rangle ,
\langle \lambda'_\delta , X'_\delta \colon \delta \in K' \rangle.$$
Then $\delta \in K'$ and $$\{ x_1 , \ldots , x_k \} \subset X'_{\delta} 
\prec (Q_{\lambda_\delta};\in {\mathbb P}_{{\lambda_\delta}},
A_{{\lambda_\delta}}){\rm , }$$
so that by (\ref{in-the-model}) we may choose some $x \in X'_{\delta}$ with
$$(Q_{\lambda_\delta};\in {\mathbb P}_{{\lambda_\delta}},
A_{{\lambda_\delta}}) \models \varphi(x,x_1,\ldots , x_k).$$ Let $r = q \cup \{
\ulcorner x \in {\dot X}_{\delta} \urcorner \}$.

By density, there is then some $y \in X_\delta$ such that $$(Q_{\lambda_\delta};\in {\mathbb P}_{{\lambda_\delta}},
A_{{\lambda_\delta}}) \models \varphi(y,x_1,\ldots , x_k).$$

The proof of (C.9) is similar.
Let again $\delta \in K$. Let $E \subset {\mathbb P}_{\lambda_\delta} \cap X_\delta^g$ be dense in ${\mathbb P}_{\lambda_\delta} \cap X_\delta$,
and $r \in E$ iff $r \in {\mathbb P}_{\lambda_\delta} \cap X_\delta$
and
\begin{eqnarray}\label{defn_of_E}
(Q_{\lambda_\delta};\in {\mathbb P}_{{\lambda_\delta}},
A_{{\lambda_\delta}}) \models \varphi(r,x_1, \ldots , x_k).
\end{eqnarray}
Let $p \in g$ be such that $\{ \ulcorner x_1 \in {\dot X}_\delta \urcorner , \ldots ,
\ulcorner x_k \in {\dot X}_\delta \urcorner , \ulcorner \delta \mapsto \lambda_\delta \urcorner
\} \subset p$. Let $q \leq p$, and again let
$\Sigma$ be a syntactical certificate for $q$ whose associated semantical certificate is
$${\mathfrak C}' = \langle M'_i , \pi'_{ij} , N'_i , \sigma'_{ij} \colon i \leq j \leq \omega_1 \rangle , 
\langle (k'_n , \alpha'_n ) \colon n<\omega \rangle ,
\langle \lambda'_\delta , X'_\delta \colon \delta \in K' \rangle.$$
Then $[\Sigma]^{<\omega} \cap X'_\delta$ has an element, say $r$, such that (\ref{defn_of_E}) holds true. Let $s = q \cup r \cup \{ \ulcorner r \in {\dot X}_\delta \urcorner \}$.

By density, then, $g \cap X_\delta \cap E \not= \emptyset$. \hfill $\square$

\begin{lemma}\label{P_stat_pres}
Let $g$ be ${\mathbb P}$-generic over $V$. Let 
$${\mathfrak C} = \langle M_i , \pi_{ij} , N_i , \sigma_{ij} \colon i \leq j \leq \omega_1 \rangle , 
\langle (k_n , \alpha_n ) \colon n<\omega \rangle ,
\langle \lambda_\delta , X_\delta \colon \delta \in K \rangle$$
be the semantic certificate associated with the syntactic certificate $\bigcup g$.
Let 
$$N_{\omega_1} = (N_{\omega_1};\in,A,I^*){\rm , }$$ and let
$T \in ({\cal P}(\omega_1) \cap N_{\omega_1}) \setminus I^*$.
Then $T$ is stationary in $V[g]$.
\end{lemma}

If ${\mathfrak C}$, $I^*$, and $T$ are as in the statement of Lemma \ref{P_stat_pres},
then by Lemma \ref{generics_give_certificates} and (C.3) and (C.4) we will have that
$({\sf NS}_{\omega_1})^V = I^* \cap V$, so that the conclusion of Lemma \ref{P_stat_pres} also gives that ${\mathbb P}$ preserves the stationarity of $T$. In other words:

\begin{corollary}
${\mathbb P}$ preserves stationary subsets of $\omega_1$.
\end{corollary}

{\sc Proof} of Lemma \ref{P_stat_pres}. Let ${\dot N}_{\omega_1} \in
V^{\mathbb P}$
be a canonical name for $N_{\omega_1}$, and let ${\dot I}^* \in
V^{\mathbb P}$
be a canonical name for $I^*$.
Let ${\bar p} \in g$, ${\dot C}$, ${\dot S} \in V^{\mathbb P}$, and
$i<\omega_1$ and $n<\omega$ be such that
\begin{enumerate}
\item[(i)] $T = {\dot S}^g$,
\item[(ii)] ${\bar p} \Vdash$ ``${\dot C} \subset \omega_1$ is club,'' 
\item[(iii)] ${\bar p} \Vdash$ ``${\dot S} \in ({\cal P}(\omega_1) \cap {\dot N}_{\omega_1}) \setminus {\dot I}^*$,'' and 
\item[(vi)] ${\bar p} \Vdash$ ``${\dot S}$ is  
represented by $[i,{\dot n}]$ in the term model producing ${\dot N}_{\omega_1}$.'' 
\end{enumerate}
We may and shall also assume that
\begin{eqnarray}\label{right-thing}
\ulcorner {\dot N}_i \models {\dot n} \mbox{ is a subset of the first uncountable cardinal, yet } {\dot n} \notin {\dot I} \urcorner \in {\bar p}.
\end{eqnarray} 
Let $p \leq {\bar p}$ be arbitrary. We aim to produce some $q \leq p$ and some $\delta<\omega_1$ such that $q \Vdash {\check \delta} \in {\dot C} \cap {\dot S}$,
see Claim \ref{claim2.10} below. 

For $\xi<\omega_1$, let $$D_\xi = \{ q \leq p \colon \exists \eta \geq \xi \, (\eta < \omega_1 \wedge q \Vdash {\check \eta} \in {\dot C}) \}{\rm, }$$ so that $D_\xi$ is open dense below $p$. Let $$E = \{ (q,\eta) \in {\mathbb P} \times \omega_1 \colon
q \Vdash {\check \eta} \in {\dot C} \}.$$ Let us write $$\tau = ((D_\xi \colon \xi<\omega_1),E).$$
We may and shall identify $\tau$ with some subset of $H_\kappa$ which codes $\tau$.

By ($\Diamond({\mathbb P})$), we may pick some 
$\lambda \in C$ such that $p \in {\mathbb P}_\lambda$ and
\begin{eqnarray}\label{choice_of_lambda}
(Q_\lambda;\in,{\mathbb P}_\lambda,A_\lambda) \prec (H_\kappa;\in,{\mathbb P},\tau).
\end{eqnarray}

Let $h$ be ${\rm Col}(\omega,2^{\aleph_2})$-generic over $V$, and let 
$g' \in V[h]$ be a filter on ${\mathbb P}_\lambda$ such that 
$p \in g'$
and $g'$ meets every dense set which is definable over $(N_\lambda;\in,{\mathbb P}_\lambda,A_\lambda)$ from parameters in $N_\lambda$.
By Lemma \ref{generics_give_certificates}, $\bigcup g'$ is a syntactic certificate for $p$,
and we may let
$$\langle M_i' , \pi_{ij}' , N_i' , \sigma_{ij}' \colon i \leq j \leq \omega_1 \rangle , 
\langle (k_n' , \alpha_n' ) \colon n<\omega \rangle ,
\langle \lambda'_\delta , X_\delta' \colon \delta \in K' \rangle$$
be the associated semantic certificate. In particular, $K' \subset \lambda$.

Let $S$ denote the subset of $\omega_1$ which is represented by $[i,{\dot n}]$ in the term model giving $N_{\omega_1}'$, so that 
if $N_{\omega_1}'=(N_{\omega_1}',\in,A,I')$, then by (\ref{right-thing}),
\begin{eqnarray}\label{S_positive}
S \in ({\cal P}(\omega_1) \cap N'_{\omega_1})
\setminus I'.
\end{eqnarray}
Let us also write $\rho=\omega_1^{V[h]} = (2^{\aleph_2})^{+V}$. 
Inside $V[h]$, we may extend $\langle N'_i , \sigma'_{ij} \colon i \leq j \leq \omega_1 \rangle$ to a generic iteration $$\langle N'_i , \sigma'_{ij} \colon i \leq j \leq \rho \rangle$$ such that 
\begin{eqnarray}
\omega_1 \in \sigma'_{\omega_1,\omega_1+1}(S).
\end{eqnarray}
This is possible as $\omega_1^{N'_{\omega_1}} = {\rm sup} \{ \omega_1^{N_j} \colon j<\omega_1 \} = \omega_1$ and by (\ref{S_positive}).
Let $$\langle M'_i , \pi'_{ij} \colon i \leq j \leq \rho \rangle = \sigma_{0,\rho}(\langle M'_i , \pi'_{ij} \colon i \leq j \leq \omega_1^{N'_0} \rangle){\rm , }$$ so that $\langle M'_i , \pi'_{ij} \colon i \leq j \leq \rho \rangle$ is an extension of $\langle M'_i , \pi'_{ij} \colon i \leq j \leq \omega_1 \rangle$.

Recalling (\ref{Momega1}), we may lift $\langle M'_i , \pi'_{ij} \colon \omega_1 \leq i \leq j \leq \rho \rangle$ to a generic iteration $$\langle M_i^+ , \pi_{ij}^+ \colon \omega_1 \leq i \leq j \leq \rho \rangle$$ of $V$. Let us write $M=M_\rho^+$ and $\pi=\pi_{\omega_1,\rho}^+$. 

The key point is now that $\langle M_i' , \pi'_{ij} , N'_i , \sigma'_{ij} \colon i \leq j \leq \rho \rangle$ may be used to extend $\pi \mbox{''} \bigcup g'$ to a
syntactic certificate 
\begin{eqnarray}\label{supset}
\Sigma \supset \pi \mbox{''} \bigcup g'
\end{eqnarray} 
for $\pi(p)$ 
in the followig manner.
Let 
$K^* = K' \cup \{ \omega_1 \}$. For $\delta \in K'$, let $\lambda_\delta^* =
\pi(\lambda_\delta')$ and $X_\delta^* =
\pi \mbox{''} X_\delta'$. Also, write $\lambda_{\omega_1}^* = \pi(\lambda)$ and $X_{\omega_1}^* = \pi \mbox{''} Q_\lambda$. Let 
$${\mathfrak C}^* =
\langle M'_i , \pi'_{ij} , N'_i , \sigma'_{ij} \colon i \leq j \leq \rho \rangle \mbox{, } 
\langle (k'_n , \pi(\alpha'_n) ) \colon n<\omega \rangle  \mbox{, } 
\langle \lambda_\delta^* , X_\delta^* \colon \delta \in K^* \rangle.$$
It is then straightforward to verify that ${\mathfrak C}^*$ is a semantic certificate
for $\pi(p)$, and that in fact there is some syntactic certificate $\Sigma$ as in
(\ref{supset}) such that ${\mathfrak C}^*$ is certified by $\Sigma$.

Now let $[{\dot m}]_{\omega_1+1}$ represent $\sigma'_{\omega_1 \omega_1+1}(S)$ in the term model
for $N_{\omega_1+1}'$ provided by $\Sigma$, so that\footnote{Here, ${\dot \sigma}_{i \omega_1+1}$ and ${\dot N}_{\omega_1+1}$ are terms of the language associated with $\pi({\mathbb P}_\lambda)$ and $\ulcorner {\dot \sigma}_{i \omega_1+1}({\dot n}) = {\dot m} \urcorner$ and $
\ulcorner {\dot N}_{\omega_1+1} \models \omega_1 \in {\dot m} \urcorner$ are formulae of that language.}  
$$\{ \ulcorner {\dot \sigma}_{i \omega_1+1}({\dot n}) = {\dot m} \urcorner ,
\ulcorner {\dot N}_{\omega_1+1} \models \omega_1 \in {\dot m} \urcorner \}
\subset \Sigma{\rm , }$$
in other words,
\begin{eqnarray}\label{half_of_new_p}
\pi(p) \cup \{ \ulcorner {\dot \sigma}_{i \omega_1+1}({\dot n}) = {\dot m} \urcorner ,
\ulcorner {\dot N}_{\omega_1+1} \models \omega_1 \in {\dot m} \urcorner \}
\mbox{ is certified by } \Sigma .
\end{eqnarray}

Let us now define 
\begin{eqnarray}\label{defn_q}
q^* = \pi(p) \cup \{
\ulcorner {\dot \sigma}_{i \omega_1+1}({\dot n}) = {\dot m} \urcorner ,
\ulcorner {\dot N}_{\omega_1+1} \models \omega_1 \in {\dot m} \urcorner ,
\ulcorner \omega_1 \mapsto \pi(\lambda)  \urcorner \}.
\end{eqnarray}

We thus established the following.

\begin{claim}\label{claim_1}
$q^* \in \pi({\mathbb P})$, as being certified by $\Sigma$.
\end{claim}

The elementarity of
$\pi \colon V \rightarrow M_\rho^+$ then gives some $\delta<\omega_1$ and some $\mu<\kappa$
such that
\begin{eqnarray}
q = p \cup \{ \ulcorner {\dot \sigma}_{i \delta+1}({\dot n}) = {\dot m} \urcorner ,
\ulcorner {\dot N}_{\delta+1} \models \delta \in {\dot m} \urcorner , \ulcorner \delta \mapsto \lambda \urcorner \} \in {\mathbb P}.
\end{eqnarray}

\begin{claim}\label{claim_2}\label{claim2.10}
$q \Vdash {\check \delta} \in {\dot C} \cap {\dot S}$.
\end{claim}

{\sc Proof} of Claim \ref{claim_2}. $q \Vdash {\check \delta} \in {\dot S}$ readily follows from $\{ \ulcorner {\dot \sigma}_{i \delta+1}({\dot n}) = {\dot m} \urcorner ,
\ulcorner {\dot N}_{\delta+1} \models \delta \in {\dot m} \urcorner \} \subset q$, the fact that ${\bar p} \geq p$ forces that ${\dot S}$ is 
represented by $[i,{\dot n}]$ in the term model giving ${\dot N}_{\omega_1}$, and the fact that
by Claim \ref{claim2.4}, $[\delta]_{\delta+1}$ represents $\delta$ in the model $N_{\delta+1}$
of any semantic certificate for $q$. 

Let us now show that $q \Vdash {\check \delta} \in {\dot C}$. 
We will in fact show that $q$ forces that ${\check \delta}$ is a limit point of ${\dot C}$.
Otherwise
there is some $r \leq q$ and some $\eta < \delta$ such that 
\begin{eqnarray}\label{21}
r \Vdash {\dot C} \cap {\check \delta} \subset {\check \eta}.
\end{eqnarray}
Let 
\begin{eqnarray}\label{prime}
\langle M'_i , \pi'_{ij} , N'_i , \sigma'_{ij} \colon i \leq j \leq \omega_1 \rangle , 
\langle (k'_n , \alpha'_n ) \colon n<\omega \rangle ,
\langle \lambda'_{\bar \delta} , X'_{\bar \delta} \colon {\bar \delta} \in K' \rangle
\end{eqnarray}
certify $r$. We must have that 
\begin{enumerate}
\item[(a)] $\delta \in K'$, 
\item[(b)] $X'_\delta \prec (Q_\lambda; \in , {\mathbb P}_\lambda , A_\lambda)$,
\item[(c)] $X'_\delta \cap \omega_1 = \delta$,
and
\item[(d)] for some $\Sigma$ such that the objects from (\ref{prime}) are certified by $\Sigma$, $[\Sigma]^{<\omega} \cap X'_\delta \cap
E \not= \emptyset$ for every $E \subset {\mathbb P}_{\lambda}$ which is dense in ${\mathbb P}_{\lambda} \cap 
X'_\delta$ and definable over the structure 
$$(Q_\lambda;\in, {\mathbb P}_\lambda , A_\lambda)$$
from parameters in $X'_\delta$.
\end{enumerate}
Notice that $A_\lambda = \tau \cap Q_\lambda$, and hence $A_\lambda$ may be identified 
with $((D_\xi \cap Q_\lambda \colon \xi < \omega_1), E \cap Q_\lambda )$.
As $\eta < \delta \subset X_\delta'$, $D_\eta$ is definable over
the structure $$(Q_\lambda;\in, {\mathbb P}_\lambda , A_\lambda)$$
from a parameter in $X'_\delta$. By
(\ref{choice_of_lambda}), $D_\eta \cap Q_\lambda$ is dense in ${\mathbb P}_\lambda$.
By (d) above, there is then some
$s \in [\Sigma]^{<\omega} \cap X'_\delta \cap D_\eta \cap Q_\lambda$. 

By (\ref{choice_of_lambda}) again, the unique smallest $\eta' \geq \eta$ with $s \Vdash {\check \eta}' \in {\dot C}$ must be in $X'_\delta$, hence $\eta' < \delta$ by (c) above. 
By Lemma \ref{+++}, $s$ is compatible with $r$.
We have reached a contradiction with (\ref{21}).
\hfill $\square$

\end{document}